Your roommate is working on his bicycle and has the bikeupside down. He spins the 62.0 cm-diameter wheel, and you notice that a pebble stuck in thetread goes by three times every second.
What is the pebble's speed? and what is thepebbles acceleration? are the two questions I must answer. I haveno idea how to set up this problem. Please help.
Thanks.
Answer #1
Please remember to rate!
--
If the pebble goes by 3 times per second, then the angularspeed is given in radian by:
--
3 rev/s*2π = 18.84 rad/s
--
Tangential speed is found with:
--
v = rω
v = (0.62/2)18.84 rad/s
v = 5.84 m/s
--
--
--
Acceleration is found with:
--
a = v^2/r
a = 5.84^2/(0.62/2)
a = 110 m/s^2
--
--
Hope this helps.